# MGDA Converges under Generalized Smoothness, Provably

### v4

@@ D.2 Formal Version and Its Proof of Theorem 4
We then have the following convergence rate for Algorithm 4.
@@
Suppose Assumptions 1 and 2 are satisfied, and we choose constant step sizes that $\beta\leq\frac{\epsilon^{2}}{C_{1}^{\prime 2}},\alpha\leq\min\left(c_{1}\beta,\sqrt{\frac{2c_{3}^{\prime}}{K\ell(M+1)M^{2}T}},\frac{2c_{4}^{\prime}}{\beta C_{1}^{\prime 2}T},\frac{\epsilon^{2}}{K\ell(M+1)M^{2}}\right),\rho\leq\left(\frac{\epsilon^{2}}{2},\frac{c_{2}^{\prime}}{\alpha T}\right)$, and $T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)$. We have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}$
@@
Following similar steps in Section C.1, we also prove that for any $i\in K$ and $t\leq T$, we have that $f_{i}(x_{t})-f_{i}^{*}\leq F$ by induction.
@@
Induction step: assume that for any $i\in[K]$ and $t\leq k<T$, $f_{i}(x_{t})-f_{i}^{*}\leq F$ holds. We then prove $f_{i}(x_{k+1})-f_{i}^{*}\leq F$ holds for any $i\in[K]$. Following similar steps in Section C.1, we have
@@
where $(i)$ follows from Cauchy-Schwartz inequality and $C_{1}^{\prime}=\sqrt{K}M^{2}+\rho+\frac{\alpha^{2}K\ell(M+1)M^{2}}{2}$, and $(ii)$ follows from Lemma 5. Then we have
@@
Then substitute the above inequality into eq. 33, we can obtain
@@
since we have that $\frac{\alpha}{\beta}\leq c_{1}$, $\alpha\rho T\leq c_{2}^{\prime}$, $\frac{\alpha^{2}K\ell(M+1)M^{2}T}{2}\leq c_{3}\prime$, $\frac{\alpha\beta(C_{1}^{\prime})^{2}T}{2}\leq c_{4}^{\prime}$. Now we finish the induction step and can show that $f_{i}(x_{k})-f_{i}^{*}\leq F$ and section D.2 hold for all $k<T$ and $i\in[K]$. Specifically, for $\alpha\leq\frac{1}{\ell(M+1)}$, we have
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}\leq\epsilon^{2}.$
@@
D.3 Formal Version and Its Proof of Theorem 5
@@
Suppose Assumptions 1 and 2 are satisfied. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},N\sim\mathcal{O}(\epsilon^{-2}),\beta\leq\min\left(\frac{\epsilon^{2}\rho}{(C_{1}^{\prime})^{2}},\epsilon^{2}\right),\alpha\leq\min\left(c_{1}\beta,\frac{2c_{3}^{\prime}}{\beta c_{1}^{\prime}T},\frac{1}{\ell(M+1)},\frac{\beta\rho\epsilon}{2L_{w}\sqrt{M}},\frac{\rho\epsilon^{2}}{2L_{w}MC_{1}^{\prime}}\right),T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)\sim\Theta(\epsilon^{-11})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\frac{c_{2}^{\prime}}{2T\alpha}\right)\sim\mathcal{O}(\epsilon^{2})$. The CA distance in every iteration takes the order of $\mathcal{O}(\epsilon)$.
@@
$-2\frac{\beta}{\alpha}\langle R(x_{t}),w_{t}-w_{t,\rho}^{*}\rangle+\beta^{2}\|\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+R(x_{t})+\rho w_{t}\|^{2}$
$\overset{(ii)}{\leq}$ $(1-2\beta\rho)\|w_{t}-w_{t,\rho}^{*}\|^{2}+2\frac{\beta}{\alpha}\|R(x_{t})\|+\beta^{2}(\rho+\sqrt{K}M^{2}+\|R(x_{t})\|)^{2},$ (37)
@@
$\leq$ $2\alpha\beta L_{w}M(\sqrt{K}M^{2}+\|R(x_{t})\|+\rho)+\beta\rho\|w_{t}-w_{t,\rho}^{*}\|^{2}+\frac{4\alpha^{2}L_{w}^{2}M}{\beta\rho}.$ (39)
Then substituting section D.3, eq. 38 and section D.3 into eq. 36, we have
@@
$+\alpha^{2}L_{w}^{2}M+2\alpha\beta L_{w}M(\sqrt{K}M^{2}+\|R(x_{t})\|+\rho)+\frac{4\alpha^{2}L_{w}^{2}M}{\beta\rho}$
@@
$+\alpha^{2}L_{w}^{2}M+2\alpha\beta L_{w}MC_{1}^{\prime}+\frac{4\alpha^{2}L_{w}^{2}M}{\beta\rho},$
where the last inequality follows from Lemma 5 and $C_{1}^{\prime}=\sqrt{K}M^{2}+\rho+\frac{\alpha^{2}K\ell(M+1)M^{2}}{2}$. Then we do telescoping over $t=0,1,...,T-1$
@@
$+\frac{\alpha^{2}}{\beta\rho}L_{w}^{2}M+\frac{2\alpha L_{w}M}{\rho}C_{1}^{\prime}+\frac{4\alpha^{2}L_{w}^{2}M}{\beta^{2}\rho^{2}}.$
Then substituting the above inequality into section D.3, we have
@@
Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 4, according to Theorem 10.5 [14], we have
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|=\mathcal{O}(\epsilon)$ as $N\sim\mathcal{O}(\rho^{-1})$. CA distance takes the order of $\epsilon$ in every iteration by choosing $\rho\sim\mathcal{O}(\epsilon^{2}),\beta\sim\mathcal{O}(\epsilon^{4})$, $\alpha\sim\mathcal{O}(\epsilon^{9})$, and $N\sim\mathcal{O}(\epsilon^{-2})$. The proof is complete. ∎
D.4 Formal Version of Its Proof of Theorem 6
Let $\alpha,\beta,\rho,T$ satisfy all requirements for Theorem 2 with $\delta<\frac{1}{2}$. Moreover, for $\rho\sim\mathcal{O}(\epsilon^{2}),N\sim\mathcal{O}(\epsilon^{-2}),\beta\leq\frac{\delta\rho\epsilon^{2}}{60(1+KM^{4})}\sim\mathcal{O}(\epsilon^{4}),n_{s}\geq\max\{\frac{1}{K\sigma},\frac{36K\sigma(6+20\beta\rho)}{\delta\rho^{2}\epsilon^{2}}\}\sim\mathcal{O}(\epsilon^{-6})$ and $\alpha\leq\sqrt{\frac{\delta\beta^{2}\rho^{2}\epsilon^{2}}{12L_{w}^{2}(2M^{2}+4\sigma^{2})(\beta\rho+1)}}\sim\mathcal{O}(\epsilon^{9})$ and $T\sim\Theta(\epsilon^{-11})$, we have the following theorem:
@@ D.4 Formal Version of Its Proof of Theorem 6
$\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\sim\mathcal{O}(\epsilon),$
@@
Base case: Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 1, according to Theorem 10.5 [14], we have by choosing $\beta^{\prime}\in(0,\frac{1}{M^{2}}]$
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|^{2}=\mathcal{O}(\frac{\delta}{2}\epsilon^{2})$ as $N\sim\mathcal{O}(\rho^{-1})$.
Induction: Assume we have that $\mathbb{E}[\|w_{t}-w_{t,\rho}^{*}\|^{2}|\tau=T]\leq\epsilon^{2}$, we will show that $\mathbb{E}[\|w_{t+1}-w_{t+1,\rho}^{*}\|^{2}|\tau=T]\leq\epsilon^{2}$ holds for any $t<\tau$ in the following proof. We first divide $\|w_{t+1}-w_{t+1,\rho}^{*}\|^{2}$ into three parts:
@@
where $(i)$ follows from the non-expansiveness of projection and $(ii)$ follows from properties of strong convexity and Cauchy-Schwartz inequality. Taking the conditional expectation of (D.4), we have that for any $a_{1}>0$,
@@
$+\frac{\delta}{2}(1-2\beta\rho)\epsilon^{2}+5\beta^{2}\rho^{2}+5\beta^{2}KM^{4}+5\beta^{2}\mathbb{E}[M\|\epsilon_{t,2}\|^{2}|\tau=T]$
$+5\beta^{2}\mathbb{E}[M\|\epsilon_{t,3}\|^{2}|\tau=T]+5\beta^{2}\mathbb{E}[\|\epsilon_{t,2}\|^{2}\|\epsilon_{t,3}\|^{2}|\tau=T],$ (43)
@@
where the first inequality is due to Lemma 3, where $L_{w}\sim\mathcal{O}(\rho)$. Then for the last term on the RHS in eq. 41, for any $a_{2}>0,a_{3}>0$, we have that
@@
$\leq$ $a_{2}(5\beta^{2}\rho^{2}+5\beta^{2}KM^{4}+5\beta^{2}\mathbb{E}[M\|\epsilon_{t,2}\|^{2}|\tau=T]$
$+5\beta^{2}\mathbb{E}[M\|\epsilon_{t,3}\|^{2}|\tau=T]+5\beta^{2}\mathbb{E}[\|\epsilon_{t,2}\|^{2}|\|\epsilon_{t,3}\|^{2}|\tau=T])$
@@
where $(i)$ follows from the non-expansiveness of projection and (D.4), and the last inequality is from (D.4). Then substituting section D.4, section D.4 and section D.4 into eq. 41, we have
@@
$+M\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\mathbb{E}[\|\varepsilon_{t,2}\|^{2}|\tau=T]$
$+M\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\mathbb{E}[\|\varepsilon_{t,3}\|^{2}|\tau=T]$
@@
$+M\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\left({\frac{4K\sigma}{n_{s}}}+\frac{{2}K^{2}\sigma^{2}}{n_{s}^{2}}\right)$
$+\left(1+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\alpha^{2}L_{w}^{2}\left(2M^{2}+\frac{4\sigma^{2}}{n_{s}}\right),$ (46)
@@
$\leq\frac{\sqrt{2}K\sigma^{2}}{n_{s}}\sigma.$
According to (D.4), with $a_{1}=0.5\rho,a_{2}=1,a_{3}=0.5\beta\rho,\beta\leq\frac{\delta\rho\epsilon^{2}}{60(1+KM^{4})},n_{s}\geq\max\{\frac{1}{K\sigma},\frac{36K\sigma(6+20\beta\rho)}{\delta\rho^{2}\epsilon^{2}}\}$ and $\alpha\leq\sqrt{\frac{\delta\beta^{2}\rho^{2}\epsilon^{2}}{12L_{w}^{2}(2M^{2}+4\sigma^{2})(\beta\rho+1)}}$, we have that
@@
where the last inequality is because our parameters satisfy all the requirements in Theorem 2, thus $\mathbb{P}(\tau<T)\leq\frac{\delta}{2}$. Then based on (D.4), by setting $\rho\sim\mathcal{O}(\epsilon^{2})$, we have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|\sim\mathcal{O}(\epsilon)$ with probability at least $1-\delta$ for each iteration $t$, which completes the proof. ∎

### v5

@@ E.2 Formal Version and Its Proof of Theorem 4
and $C_{1}^{\prime}\geq\sqrt{K}M^{2}+\rho+\frac{\alpha K\ell(M+1)M^{2}}{2}$. We then have the following convergence rate for Algorithm 4.
@@
Suppose Assumptions 1 and 2 are satisfied, and we choose constant step sizes that $\beta\leq\frac{\epsilon^{2}}{C_{1}^{\prime 2}},\alpha\leq\min\left(c_{1}\beta,\sqrt{\frac{2c_{3}^{\prime}}{K\ell(M+1)M^{2}T}},\frac{2c_{4}^{\prime}}{\beta C_{1}^{\prime 2}T},\frac{\epsilon^{2}}{K\ell(M+1)M^{2}}\right),\rho\leq\min\left(\frac{\epsilon^{2}}{2},\frac{c_{2}^{\prime}}{\alpha T}\right)$, and $T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)$. We have
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}).$
@@
Following similar steps in Section D.1, we also prove that for any $i\in K$ and $t\leq T$, we have that $f_{i}(x_{t})-f_{i}^{*}\leq F$ by induction.
@@
Induction step: assume that for any $i\in[K]$ and $t\leq k<T$, $f_{i}(x_{t})-f_{i}^{*}\leq F$ holds. We then prove $f_{i}(x_{k+1})-f_{i}^{*}\leq F$ holds for any $i\in[K]$. Following similar steps in Section D.1, we have
@@
where $(i)$ follows from Cauchy-Schwartz inequality and $C_{1}^{\prime}\geq\sqrt{K}M^{2}+\rho+\frac{\alpha K\ell(M+1)M^{2}}{2}$, and $(ii)$ follows from Lemma 5. Then we have
@@
Then substituting the above inequality into eq. 32, we can obtain
@@
since we have that $\frac{\alpha}{\beta}\leq c_{1}$, $\alpha\rho T\leq c_{2}^{\prime}$, $\frac{\alpha^{2}K\ell(M+1)M^{2}T}{2}\leq c_{3}\prime$, $\frac{\alpha\beta(C_{1}^{\prime})^{2}T}{2}\leq c_{4}^{\prime}$. Now we finish the induction step and can show that $f_{i}(x_{k})-f_{i}^{*}\leq F$ and section E.2 hold for all $k<T$ and $i\in[K]$. Specifically, for $\alpha\leq\frac{1}{\ell(M+1)}$, we have
@@
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla F(x_{t})w_{t}\|^{2}=\mathcal{O}(\epsilon^{2}).$
@@
E.3 Formal Version and Its Proof of Theorem 5
@@
Suppose Assumptions 1 and 2 are satisfied. We choose $\beta^{\prime}\leq\frac{1}{M^{2}},N=\Omega(\epsilon^{-2}),{C_{1}^{\prime}\geq\sqrt{K}M^{2}+\rho+\frac{\alpha K\ell(M+1)M^{2}}{2}},\beta\leq\min\left(\frac{\epsilon^{2}\rho}{(C_{1}^{\prime})^{2}},\epsilon^{2}\right),\alpha\leq\min\left(c_{1}\beta,\frac{2c_{3}^{\prime}}{\beta c_{1}^{\prime}T},\frac{1}{\ell(M+1)},\frac{\beta\rho\epsilon}{2L_{w}M},\frac{\rho\epsilon^{2}}{2L_{w}MC_{1}^{\prime}}\right),T\geq\max\left(\frac{10\Delta}{\alpha\epsilon^{2}},\frac{10}{\epsilon^{2}\beta}\right)=\Theta(\epsilon^{-11})$, and $\rho\leq\min\left(\frac{\epsilon^{2}}{20},\frac{c_{2}^{\prime}}{2T\alpha}\right)=\mathcal{O}(\epsilon^{2})$. The CA distance in every iteration takes the order of $\mathcal{O}(\epsilon)$.
@@
$-2\frac{\beta}{\alpha}\langle R(x_{t}),w_{t}-w_{t,\rho}^{*}\rangle+\beta^{2}\left\|\nabla F(x_{t})^{\top}\nabla F(x_{t})w_{t}+\frac{R(x_{t})}{\alpha}+\rho w_{t}\right\|^{2}$
$\overset{(ii)}{\leq}$ $(1-2\beta\rho)\|w_{t}-w_{t,\rho}^{*}\|^{2}+2\frac{\beta}{\alpha}\|R(x_{t})\|+\beta^{2}\left(\rho+\sqrt{K}M^{2}+\frac{\|R(x_{t})\|}{\alpha}\right)^{2},$ (36)
@@
$\leq$ $2\alpha\beta L_{w}M\left(\sqrt{K}M^{2}+\frac{\|R(x_{t})\|}{\alpha}+\rho\right)+\beta\rho\|w_{t}-w_{t,\rho}^{*}\|^{2}+\frac{4\alpha^{2}L_{w}^{2}M^{2}}{\beta\rho}.$ (38)
Then substituting section E.3, eq. 37 and section E.3 into eq. 35, we have
@@
$+\alpha^{2}L_{w}^{2}M+2\alpha\beta L_{w}M\left(\sqrt{K}M^{2}+\frac{\|R(x_{t})\|}{\alpha}+\rho\right)+\frac{4\alpha^{2}L_{w}^{2}M^{2}}{\beta\rho}$
@@
$+\alpha^{2}L_{w}^{2}M+2\alpha\beta L_{w}MC_{1}^{\prime}+\frac{4\alpha^{2}L_{w}^{2}M^{2}}{\beta\rho},$
where the last inequality follows from Lemma 5 and $C_{1}^{\prime}\geq\sqrt{K}M^{2}+\rho+\frac{\alpha K\ell(M+1)M^{2}}{2}$. Then we do telescoping over $t=0,1,...,T-1$
@@
$+\frac{\alpha^{2}}{\beta\rho}L_{w}^{2}M+\frac{2\alpha L_{w}M}{\rho}C_{1}^{\prime}+\frac{4\alpha^{2}L_{w}^{2}M^{2}}{\beta^{2}\rho^{2}}.$
Then substituting the above inequality into section E.3, we have
@@
Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 4, according to Theorem 10.5 (Garrigos & Gower, 2023), we have
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|=\mathcal{O}(\epsilon)$ as $N=\Omega(\rho^{-1})$. CA distance takes the order of $\epsilon$ in every iteration by choosing $\rho=\mathcal{O}(\epsilon^{2}),\beta=\mathcal{O}(\epsilon^{4})$, $\alpha=\mathcal{O}(\epsilon^{9})$, and $N=\Omega(\epsilon^{-2})$. The proof is complete. ∎
E.4 Formal Version of Its Proof of Theorem 6
Let $\alpha,\beta,\rho,T$ satisfy all requirements for Theorem 2 with $\delta<\frac{1}{2}$. Moreover, for $\rho=\mathcal{O}(\epsilon^{2}),N=\Omega(\epsilon^{-2}),\beta\leq\frac{\delta\rho\epsilon^{2}}{60(1+KM^{4})}=\mathcal{O}(\epsilon^{4}),n_{s}\geq\max\{{K\sigma^{2}},\frac{36K\sigma^{2}M^{2}(6+20\beta\rho)}{\delta\rho^{2}\epsilon^{2}}\}=\Omega(\epsilon^{-6})$ and $\alpha\leq\sqrt{\frac{\delta\beta^{2}\rho^{2}\epsilon^{2}}{12L_{w}^{2}(2M^{2}+4K\sigma^{2})(\beta\rho+1)}}=\mathcal{O}(\epsilon^{9})$ and $T=\Theta(\epsilon^{-11})$, we have the following theorem:
@@ E.4 Formal Version of Its Proof of Theorem 6
$\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|=\mathcal{O}(\epsilon),$
@@
Base case: Since we run projected gradient descent for the strongly convex function $J(w_{n})=\frac{1}{2}\|\nabla F(x_{0})w_{n}\|^{2}+\frac{\rho}{2}\|w_{n}\|^{2}$ in the N-loop in Algorithm 1, according to Theorem 10.5 (Garrigos & Gower, 2023), we have by choosing $\beta^{\prime}\in(0,\frac{1}{M^{2}}]$
@@
Thus, $\|w_{0}-w_{0,\rho}^{*}\|^{2}=\mathcal{O}(\frac{\delta}{2}\epsilon^{2})$ as $N=\Omega(\rho^{-1})$.
Induction: Assume we have that $\mathbb{E}[\|w_{t}-w_{t,\rho}^{*}\|^{2}|\tau=T]\leq\frac{\delta}{2}\epsilon^{2}$, we will show that $\mathbb{E}[\|w_{t+1}-w_{t+1,\rho}^{*}\|^{2}|\tau=T]\leq\frac{\delta}{2}\epsilon^{2}$ holds for any $t<\tau$ in the following proof. We first divide $\|w_{t+1}-w_{t+1,\rho}^{*}\|^{2}$ into three parts:
@@
where $(i)$ follows from the non-expansiveness of projection and $(ii)$ follows from properties of strong convexity and Cauchy-Schwartz inequality. Taking the conditional expectation of equation E.4, we have that for any $a_{1}>0$,
@@
$+\frac{\delta}{2}(1-2\beta\rho)\epsilon^{2}+5\beta^{2}\rho^{2}+5\beta^{2}KM^{4}+5\beta^{2}\mathbb{E}[M^{2}\|\epsilon_{t,2}\|^{2}|\tau=T]$
$+5\beta^{2}\mathbb{E}[M^{2}\|\epsilon_{t,3}\|^{2}|\tau=T]+5\beta^{2}\mathbb{E}[\|\epsilon_{t,2}\|^{2}\|\epsilon_{t,3}\|^{2}|\tau=T],$ (42)
@@
where the first inequality is due to Lemma 3, where $L_{w}=\mathcal{O}(\rho^{-1})$. Then for the last term on the RHS in eq. 40, for any $a_{2}>0,a_{3}>0$, we have that
@@
$\leq$ $a_{2}(5\beta^{2}\rho^{2}+5\beta^{2}KM^{4}+5\beta^{2}\mathbb{E}[M^{2}\|\epsilon_{t,2}\|^{2}|\tau=T]$
$+5\beta^{2}\mathbb{E}[M^{2}\|\epsilon_{t,3}\|^{2}|\tau=T]+5\beta^{2}\mathbb{E}[\|\epsilon_{t,2}\|^{2}|\|\epsilon_{t,3}\|^{2}|\tau=T])$
@@
where $(i)$ follows from the non-expansiveness of projection and equation E.4, and the last inequality is from equation E.4. Then substituting section E.4, section E.4 and section E.4 into eq. 40, we have
@@
$+M^{2}\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\mathbb{E}[\|\varepsilon_{t,2}\|^{2}|\tau=T]$
$+M^{2}\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\mathbb{E}[\|\varepsilon_{t,3}\|^{2}|\tau=T]$
@@
$+M^{2}\left(\frac{3\beta}{a_{1}}+5\beta^{2}+5\beta^{2}a_{2}\right)\left({\frac{4K\sigma^{2}}{n_{s}}}+\frac{{2}K^{2}\sigma^{4}}{n_{s}^{2}}\right)$
$+\left(1+\frac{1}{a_{2}}+\frac{1}{a_{3}}\right)\alpha^{2}L_{w}^{2}\left(2M^{2}+\frac{4K\sigma^{2}}{n_{s}}\right),$ (45)
@@
$\leq\frac{\sqrt{2}K\sigma^{2}}{n_{s}}.$
According to equation E.4, with $a_{1}=0.5\rho,a_{2}=1,a_{3}=0.5\beta\rho,\beta\leq\frac{\delta\rho\epsilon^{2}}{60(1+KM^{4})},n_{s}\geq\max\{{K\sigma^{2}},\frac{36K\sigma^{2}M^{2}(6+20\beta\rho)}{\delta\rho^{2}\epsilon^{2}}\}$ and $\alpha\leq\sqrt{\frac{\delta\beta^{2}\rho^{2}\epsilon^{2}}{12L_{w}^{2}(2M^{2}+4K\sigma^{2})(\beta\rho+1)}}$, we have that
@@
where the last inequality is because our parameters satisfy all the requirements in Theorem 2, thus $\mathbb{P}(\tau<T)\leq\frac{\delta}{2}$. Then based on equation E.4, by setting $\rho=\mathcal{O}(\epsilon^{2})$, we have that $\|\nabla F(x_{t})w_{t}-\nabla F(x_{t})w_{t}^{*}\|=\mathcal{O}(\epsilon)$ with probability at least $1-\delta$ for each iteration $t$, which completes the proof. ∎